# Towards Agentic Intelligence for Materials Science

### v1

@@ 4.2.2 Scientific Critical Thinking
Unlike general-domain LLM reasoning, chemical reasoning must deal with uncertainty, confounded and scarce datasets, and slow, costly experimental feedback353, 367. Literature data is often noisy or conflicting, and learning frequently depends on imperfect proxy rewards from simulations, since reliable signals from laboratory experiments are scarce and costly353. At the same time, molecular discovery is inherently multi-objective, where improving one function can come at the cost of damaging another165.
Human cognition is often described as two modes of thought: System 1 (“fast thinking”) and System 2 (“slow thinking”) 175. System 2 is crucial for complex problems that require deeper and deliberate decision-making. Recent work on Energy‑Based Transformers frames “thinking” as an optimization that not only feeds richer context but also assesses the plausibility of its own predictions and dynamically determines when to halt reasoning129. Future research should aim to develop a System 2–style thinking agent to anticipate the functions and interactions of generated molecules in future physical states and encode these imagined outcomes as additional input, enabling deliberate reasoning and hypothesis generation 129. The model should also generate and evaluate alternative hypotheses, fill in missing contextual information, and connect disparate findings420. Rather than trusting all sources equally, it should assign weights to claims, hypothesize negative results, and actively resolve conflicting evidence420. To improve its causal understanding, the model should simulate counterfactual alternative molecular designs and their expected outcomes299, prioritize underexplored regions of chemical space through active learning, especially those characterized by conflicting evidence or high uncertainty in counterfactual simulations165.
@@ 4.2.3 Experiment and Simulation Planning (Decision-Making Under Uncertainty)
Another important skill for a Scientist AI is being able to plan experiments and simulations when things are uncertain353, 367. In this situation, planning is choosing the next actions—new measurements, simulations, or adjustments to the process—that are thought to be the most useful or informative based on the present hypotheses, available data, and priors93. Experiment and simulation planning, on the other hand, needs decisions to be made in order with restricted budgets, different costs, and safety or feasibility limits on which activities can be taken347.
@@
Modern materials informatics has created a wide range of tools for flexible experimental design353. Active learning and Bayesian optimization frameworks utilize uncertainty-aware surrogate models and acquisition functions to dynamically select fresh measures that either optimize information gain or enhance a specified property 367. Optimal experimental design formalizes the challenge as a trade-off between elucidating the response surface and differentiating among competing models, frequently within the confines of limited resources93, 347. These methods already allow practitioners to decide which composition, processing condition, or microstructure to evaluate next based on a current surrogate and its uncertainty estimates. However, they typically optimize scalar objectives or predictive uncertainty rather than explicitly targeting hypothesis discrimination or mechanistic insight367.
@@
In self-driving laboratory (SDL) architectures, planning experiments and automating the lab have been closely linked. Closed-loop platforms for thin-film and semiconductor discovery combine robotic synthesis and characterization with BO-based planners to autonomously search process spaces and improve multi-step recipes200. Multi-objective and restricted Bayesian optimization has facilitated these platforms in managing trade-offs among competing performance measures while adhering to intricate experimental limitations, thus enhancing Pareto fronts for conductivity, transparency, or mechanical properties 233. Recent reviews combine these changes and show that SDLs can be used as a general framework for planning closed-loop experiments in chemistry and materials science 353. From a pipeline-centric viewpoint, SDLs implement end-to-end credit assignment by directly linking model updates and planner behavior to experimental outcomes, as opposed to relying exclusively on offline proxy metrics 353.
@@
Emerging LLM-based agents enhance experiment and simulation planning by integrating language models with laboratory control systems, computational tools, and data services. ChemCrow54 and other chemistry agents show that LLMs with special tools can design reaction sequences, call simulators, query databases, and suggest follow-up experiments based on intermediate results. Autonomous research frameworks utilizing general-purpose LLMs47 have demonstrated the capacity to strategize and implement multi-step experimental campaigns, encompassing the selection of conditions, coordination of robotic platforms, and iterative enhancement of protocols 353. Recent research assesses LLM agents for the automation of particular experimental modalities, like atomic force microscopy, underscoring both the potential and the existing fragility of language-model-driven laboratory control237. These systems are starting to integrate symbolic thinking, protocol generation, and tool invocation into a single planning cycle.
@@
Even with these improvements, today’s experiment and simulation planners still fall short of expert scientists in multiple dimensions. Most Bayesian optimization and active learning frameworks focus on optimizing scalar objectives or surrogate uncertainty; they do not explicitly represent or update structured hypotheses93. Acquisition functions are usually defined over design variables instead of explanatory models or mechanisms. Consequently, campaigns seldom aim to differentiate between rival mechanistic hypotheses, delineate phase boundaries with measurable confidence, or conduct stress tests in areas of theoretical ambiguity. Ad hoc filtering is typically used to deal with practical problems like scheduling, equipment failures, maintenance windows, or safety limitations, instead of using a systematic decision-theoretic formulation 353. LLM-based planners can create flexible protocols and coordinate tools, but they are still restricted by hallucinations, poor calibration of uncertainty47, and inadequate guarantees that they will follow safety or physical rules54. Achieving scientist-level experiment and simulation planning will probably necessitate the amalgamation of uncertainty-aware optimization, explicit hypothesis representations, constraint-aware scheduling, and verifiable control of automated platforms into a cohesive framework for comprehensive materials discovery.
@@ 4.2.4 Interpretation, Explanation, and Hypothesis Revision
Within the Scientist AI system, the interpretive part of the reasoning process is what pits new data against old hypotheses and comes up with new ones that better explain the facts. Section 3.6 discussed how to make AI explainable in materials science in general. It introduced physical validity, faithfulness, and stability as ways to evaluate explanations 427. In this context, these axes measure how well explanations help revise hypotheses instead of just checking models: a Scientist AI must use explanations to figure out which hypotheses are still valid, which are clearly wrong, and where more data would best improve the set of hypotheses 269 303.
@@
Post hoc explanation techniques, such as feature attribution (e.g., LIME 300, SHAP 230) and saliency analyses, are now commonly utilized in composition-based models, graph neural networks, and other architectures 427. In materials science, these methods show which structural, compositional, or processing features affect individual predictions and whether models seem to be based on chemically plausible patterns 427 269. Attribution methods for graph-based models that pinpoint importance to atoms, bonds, or subgraphs further link predictions to specific motifs and have been evaluated on tasks with known ground-truth functional groups or structural units 9. From the standpoint of hypothesis revision, these tools facilitate a constrained form of testing: novel data can be evaluated for coherence between emphasized features and the mechanistic narrative embedded in a specific model class; however, the explanations themselves typically do not preserve or revise an explicit hypothesis set.
@@
As discussed in Sec. 3.6, sparse and closed-form models like SISSO and symbolic regression are unique in the explainability landscape 268 378. Within the framework of Scientist AI, their primary contribution is to furnish explicit candidate hypotheses as human-readable descriptors or analytical laws that are subject to direct examination, scrutiny, and modification. Descriptor-discovery workflows that incorporate SISSO-style algorithms or symbolic regression into active learning loops for electrocatalysis and other applications exemplify this potential: the system not only modifies parameters within a static model but can also transition between different functional forms as new data emerges 253. AI Feynman and other general equation-discovery frameworks show that it is possible to get governing equations from data in idealized settings359. This suggests a way to make reasoning modules that work with and choose between symbolic hypotheses instead of just updating neural weights that are hard to understand.
@@
Even though these systems have improved, they still don’t fully meet the needs of Scientist AI when it comes to interpreting and revising hypotheses. Most methods discussed in Sec. 3.6 mainly focus on local or model-centric explainability, and their faithfulness and stability when the distribution shifts or the dataset is biased are still limited 303, 427. Symbolic techniques are getting closer to clear, testable hypotheses, but they still only work in low-dimensional environments and need a lot of expert input on feature spaces and operator sets 378, 268. Emerging LLM- and agent-based systems can generate textual rationales, alternative mechanisms, or literature-grounded critiques in response to new data 231, 50, yet these explanations are seldom tied to a formal representation of a hypothesis set, nor evaluated on their ability to drive correct revisions of that set over multiple iterations of data collection. To reach scientist-level interpretation and hypothesis revision, it is therefore essential to move beyond the primarily diagnostic explainability criteria and methods of Sec. 3.6 toward explicit hypothesis tracking, uncertainty-aware selection among competing mechanistic models, and benchmarks that directly measure how effectively an AI system revises and refines scientific hypotheses within a closed-loop discovery process.
@@ 4.3 Human–AI Collaboration in Scientific Workflows
Existing molecular AI systems remain static learners trained on historical literature data and often fail to adapt to new experimental evidence or improve through real-world feedback. Specifically, future work should establish a closed-loop digital–physical workflow where a high-throughput, autonomous laboratory can synthesize AI-selected molecules with minimal human involvement38, 348. Recent work 102 proposed to position mCLM as a core agent of a human-in-the-loop autonomous laboratory, iterating through reasoning, proposal, synthesis, testing, and feedback from real-world synthesis and testing machines to enable never-ending improvement alongside human scientists, by leveraging LLM agents’ self-improvement 380, 6 and self-evolving 118 capabilities.
@@
As artificial intelligence systems mature from tools of prediction to agents of discovery, a central question has increasingly been posed: how should humans and machines collaborate in the pursuit of scientific progress 360? The further improvement of AI4MatSci will not be achieved only by algorithmic and computational power. It will also depend on the design of human–AI workflows that involve human intuition and domain insight alongside machine-scale reasoning, data integration, and hypothesis generation 149, 320. This convergence motivates the design and use of co-creative intelligence 326, in which AI could serve as a scientific collaborator.
@@
Early examples of such collaboration have already affected the materials design cycle. In prediction and generation tasks, researchers use interpretable graph neural networks (GNNs) such as ALIGNN 79 or CrabNet 369 to provide explanations for learned structure–property relationships, facilitating scientific validation and hypothesis refinement. These models support a two-way interaction between human insight and model-driven suggestions. AI can highlight correlations that are hard to detect intuitively in high-dimensional feature spaces while human scientists can test and interpret their physical plausibility. Studies in explainable materials modeling show that transparent model architectures, such as attention mechanisms revealing atomic contributions or saliency mapping over crystal graphs, can perform scientific structural reasoning from numerical output. 149. Such methods are critical for enabling trust calibration, where researchers learn not to over-trusting and under-trusting 48.
@@
Large language models (LLMs) further facilitate the interaction between AI and humans by introducing natural-language interfaces into the workflows. LLMs that are fine-tuned on domain-specific training resources such as experimental protocols and computational reports can effectively parse, summarize, and contextualize scientific results. Recent systems like MatGPT 379 are capable of solving materials language processing tasks by conducting workflows, querying crystallographic databases, and helping experimental setups by giving suggestions in natural language 76. By integrating multimodal reasoning that combines textual, numerical, and structural embeddings, LLMs can also serve as adaptive documentation and knowledge retrieval systems. These systems can evolve with laboratory data streams. In this manner, human–AI collaboration is expansive, from mere task execution to the whole scientific lifecycle: from idea creation to experiment construction and execution, validation, interpretation, and publication 226.
@@
AI systems are becoming more and more autonomous, which makes human scientists become less directly involved in low-level decision-making. This will reduce the participation of human agency and makes co-creation more difficult in some scenarios. More and more AI systems like ChemOS 302 can enhance the planning and execution of experiments in self-driving labs. In this case, human scientists should foucs more on higher-level management, ethical oversight, supervision, and conceptual innovation343. Interface design further contributes to more effective collaboration, which involves designing interpretable feedback loops between algorithmic reasoning and human comprehension 73. For instance, prospective interfaces can prevent confusion before it happens. An explainable user interface allows users to interactively explore the data and model sensitivity so they can predict what and why the AI will do. The AI decision can construct AI results as structured hypotheses, inviting human assessment to decide whether accept or deny them. These hybrid decision-making architectures align with the principles of human-in-the-loop AI, where automation and human expertise together produce more reliable outcomes than either could achieve alone 320.
@@
A second new concept at this stage is scientific collaboration networks, which were originally developed to analyze collaborations among human researchers 30 and has since broadened to include hybrid collaborations between humans and AI agents 124. In traditional scientific collaboration networks, nodes are scientists, and edges connect them if the scientists co-authored at least one paper jointly. Networks were constructed independently for separate fields for comparing the disciplinary patterns. By contrast, nodes in human-AI collaboration networks are specialized agents (software modules) along with human supervision. The framing implicitly favors human-agent teaming: humans define goals, agents create and refine hypotheses, and humans evaluate, execute, and validate results. This structure acts much like a human research team. Different members have different roles that work well together, and they talk to each other through shared interfaces to plan their work.
@@
From an epistemological point of view, the collaboration of human–AI technology signifies a paradigm shift in scientific conduct. Rather than knowledge being created solely through human deliberation, traditional science cycle by hypothesis–experiment–analysis is bounded by human thought and speed of human communication 204. Instead, co-creative workflows exist at the crossroads between symbolic reasoning on the one hand and data-driven discovery, turning this cycle into a process of feedback loop through AI 204. Transforming experimental data into organized form representations understandable to humans and machines, these systems collectively share an epistemic language that brings interpretability and reproducibility 87.
@@
Collaborative frameworks also need to address issues of attribution, authorship, and bias in the future in order to address the trust and accountability issues 259, 263. Designing new architectures and evaluation frameworks for trustworthy collaboration is becoming essential. AutoLabs 273 introduces a critic agent specifically designed to double-check the safety and feasibility, while GIFTERS 17 tries to evaluate the AI part along dimensions such as generalizable, interpretable, fair, transparent, explainable, robust, and stable trustworthiness principles. The scientific community is beginning to deal with questions like: Who owns AI-generated hypotheses? How can we peer review AI-based discoveries? Efforts such as transparent model reporting, tracking data sources, clear documentation, and continuous monitoring are essential for keeping the scientific integrity of AI-assisted research 44.Furthermore, applying fairness, transparency, and accountability principles into materials workflows is crucial for the responsible adoption of AI in both industry and academia 225.

### v2

@@ 4.2.2 Scientific Critical Thinking
Unlike general-domain LLM reasoning, chemical reasoning must deal with uncertainty, confounded and scarce datasets, and slow, costly experimental feedback355, 370. Literature data is often noisy or conflicting, and learning frequently depends on imperfect proxy rewards from simulations, since reliable signals from laboratory experiments are scarce and costly355. At the same time, molecular discovery is inherently multi-objective, where improving one function can come at the cost of damaging another166.
Human cognition is often described as two modes of thought: System 1 (“fast thinking”) and System 2 (“slow thinking”) 176. System 2 is crucial for complex problems that require deeper and deliberate decision-making. Recent work on Energy‑Based Transformers frames “thinking” as an optimization that not only feeds richer context but also assesses the plausibility of its own predictions and dynamically determines when to halt reasoning130. Future research should aim to develop a System 2–style thinking agent to anticipate the functions and interactions of generated molecules in future physical states and encode these imagined outcomes as additional input, enabling deliberate reasoning and hypothesis generation 130. The model should also generate and evaluate alternative hypotheses, fill in missing contextual information, and connect disparate findings423. Rather than trusting all sources equally, it should assign weights to claims, hypothesize negative results, and actively resolve conflicting evidence423. To improve its causal understanding, the model should simulate counterfactual alternative molecular designs and their expected outcomes301, prioritize underexplored regions of chemical space through active learning, especially those characterized by conflicting evidence or high uncertainty in counterfactual simulations166.
@@ 4.2.3 Experiment and Simulation Planning (Decision-Making Under Uncertainty)
Another important skill for a Scientist AI is being able to plan experiments and simulations when things are uncertain355, 370. In this situation, planning is choosing the next actions—new measurements, simulations, or adjustments to the process—that are thought to be the most useful or informative based on the present hypotheses, available data, and priors94. Experiment and simulation planning, on the other hand, needs decisions to be made in order with restricted budgets, different costs, and safety or feasibility limits on which activities can be taken349.
@@
Modern materials informatics has created a wide range of tools for flexible experimental design355. Active learning and Bayesian optimization frameworks utilize uncertainty-aware surrogate models and acquisition functions to dynamically select fresh measures that either optimize information gain or enhance a specified property 370. Optimal experimental design formalizes the challenge as a trade-off between elucidating the response surface and differentiating among competing models, frequently within the confines of limited resources94, 349. These methods already allow practitioners to decide which composition, processing condition, or microstructure to evaluate next based on a current surrogate and its uncertainty estimates. However, they typically optimize scalar objectives or predictive uncertainty rather than explicitly targeting hypothesis discrimination or mechanistic insight370.
@@
In self-driving laboratory (SDL) architectures, planning experiments and automating the lab have been closely linked. Closed-loop platforms for thin-film and semiconductor discovery combine robotic synthesis and characterization with BO-based planners to autonomously search process spaces and improve multi-step recipes201. Multi-objective and restricted Bayesian optimization has facilitated these platforms in managing trade-offs among competing performance measures while adhering to intricate experimental limitations, thus enhancing Pareto fronts for conductivity, transparency, or mechanical properties 235. Recent reviews combine these changes and show that SDLs can be used as a general framework for planning closed-loop experiments in chemistry and materials science 355. From a pipeline-centric viewpoint, SDLs implement end-to-end credit assignment by directly linking model updates and planner behavior to experimental outcomes, as opposed to relying exclusively on offline proxy metrics 355.
@@
Emerging LLM-based agents enhance experiment and simulation planning by integrating language models with laboratory control systems, computational tools, and data services. ChemCrow55 and other chemistry agents show that LLMs with special tools can design reaction sequences, call simulators, query databases, and suggest follow-up experiments based on intermediate results. Autonomous research frameworks utilizing general-purpose LLMs48 have demonstrated the capacity to strategize and implement multi-step experimental campaigns, encompassing the selection of conditions, coordination of robotic platforms, and iterative enhancement of protocols 355. Recent research assesses LLM agents for the automation of particular experimental modalities, like atomic force microscopy, underscoring both the potential and the existing fragility of language-model-driven laboratory control239. These systems are starting to integrate symbolic thinking, protocol generation, and tool invocation into a single planning cycle.
@@
Even with these improvements, today’s experiment and simulation planners still fall short of expert scientists in multiple dimensions. Most Bayesian optimization and active learning frameworks focus on optimizing scalar objectives or surrogate uncertainty; they do not explicitly represent or update structured hypotheses94. Acquisition functions are usually defined over design variables instead of explanatory models or mechanisms. Consequently, campaigns seldom aim to differentiate between rival mechanistic hypotheses, delineate phase boundaries with measurable confidence, or conduct stress tests in areas of theoretical ambiguity. Ad hoc filtering is typically used to deal with practical problems like scheduling, equipment failures, maintenance windows, or safety limitations, instead of using a systematic decision-theoretic formulation 355. LLM-based planners can create flexible protocols and coordinate tools, but they are still restricted by hallucinations, poor calibration of uncertainty48, and inadequate guarantees that they will follow safety or physical rules55. Achieving scientist-level experiment and simulation planning will probably necessitate the amalgamation of uncertainty-aware optimization, explicit hypothesis representations, constraint-aware scheduling, and verifiable control of automated platforms into a cohesive framework for comprehensive materials discovery.
@@ 4.2.4 Interpretation, Explanation, and Hypothesis Revision
Within the Scientist AI system, the interpretive part of the reasoning process is what pits new data against old hypotheses and comes up with new ones that better explain the facts. Section 3.6 discussed how to make AI explainable in materials science in general. It introduced physical validity, faithfulness, and stability as ways to evaluate explanations 430. In this context, these axes measure how well explanations help revise hypotheses instead of just checking models: a Scientist AI must use explanations to figure out which hypotheses are still valid, which are clearly wrong, and where more data would best improve the set of hypotheses 271 305.
@@
Post hoc explanation techniques, such as feature attribution (e.g., LIME 302, SHAP 232) and saliency analyses, are now commonly utilized in composition-based models, graph neural networks, and other architectures 430. In materials science, these methods show which structural, compositional, or processing features affect individual predictions and whether models seem to be based on chemically plausible patterns 430 271. Attribution methods for graph-based models that pinpoint importance to atoms, bonds, or subgraphs further link predictions to specific motifs and have been evaluated on tasks with known ground-truth functional groups or structural units 9. From the standpoint of hypothesis revision, these tools facilitate a constrained form of testing: novel data can be evaluated for coherence between emphasized features and the mechanistic narrative embedded in a specific model class; however, the explanations themselves typically do not preserve or revise an explicit hypothesis set.
@@
As discussed in Sec. 3.6, sparse and closed-form models like SISSO and symbolic regression are unique in the explainability landscape 270 381. Within the framework of Scientist AI, their primary contribution is to furnish explicit candidate hypotheses as human-readable descriptors or analytical laws that are subject to direct examination, scrutiny, and modification. Descriptor-discovery workflows that incorporate SISSO-style algorithms or symbolic regression into active learning loops for electrocatalysis and other applications exemplify this potential: the system not only modifies parameters within a static model but can also transition between different functional forms as new data emerges 255. AI Feynman and other general equation-discovery frameworks show that it is possible to get governing equations from data in idealized settings362. This suggests a way to make reasoning modules that work with and choose between symbolic hypotheses instead of just updating neural weights that are hard to understand.
@@
Even though these systems have improved, they still don’t fully meet the needs of Scientist AI when it comes to interpreting and revising hypotheses. Most methods discussed in Sec. 3.6 mainly focus on local or model-centric explainability, and their faithfulness and stability when the distribution shifts or the dataset is biased are still limited 305, 430. Symbolic techniques are getting closer to clear, testable hypotheses, but they still only work in low-dimensional environments and need a lot of expert input on feature spaces and operator sets 381, 270. Emerging LLM- and agent-based systems can generate textual rationales, alternative mechanisms, or literature-grounded critiques in response to new data 233, 51, yet these explanations are seldom tied to a formal representation of a hypothesis set, nor evaluated on their ability to drive correct revisions of that set over multiple iterations of data collection. To reach scientist-level interpretation and hypothesis revision, it is therefore essential to move beyond the primarily diagnostic explainability criteria and methods of Sec. 3.6 toward explicit hypothesis tracking, uncertainty-aware selection among competing mechanistic models, and benchmarks that directly measure how effectively an AI system revises and refines scientific hypotheses within a closed-loop discovery process.
@@ 4.3 Human–AI Collaboration in Scientific Workflows
Existing molecular AI systems remain static learners trained on historical literature data and often fail to adapt to new experimental evidence or improve through real-world feedback. Specifically, future work should establish a closed-loop digital–physical workflow where a high-throughput, autonomous laboratory can synthesize AI-selected molecules with minimal human involvement 39, 350. Recent work 103 proposed to position mCLM as a core agent of a human-in-the-loop autonomous laboratory, iterating through reasoning, proposal, synthesis, testing, and feedback from real-world synthesis and testing machines to enable never-ending improvement alongside human scientists, by leveraging LLM agents’ self-improvement 383, 6 and self-evolving 119 capabilities.
@@
As artificial intelligence systems mature from tools of prediction to agents of discovery, a central question has increasingly been posed: how should humans and machines collaborate in the pursuit of scientific progress 363? The further improvement of AI4MatSci will not be achieved only by algorithmic and computational power. It will also depend on the design of human–AI workflows that involve human intuition and domain insight alongside machine-scale reasoning, data integration, and hypothesis generation 150, 322. This convergence motivates the design and use of co-creative intelligence 328, in which AI could serve as a scientific collaborator.
@@
Early examples of such collaboration have already affected the materials design cycle. In prediction and generation tasks, researchers use interpretable graph neural networks (GNNs) such as ALIGNN 80 or CrabNet 372 to provide explanations for learned structure–property relationships, facilitating scientific validation and hypothesis refinement. These models support a two-way interaction between human insight and model-driven suggestions. AI can highlight correlations that are hard to detect intuitively in high-dimensional feature spaces while human scientists can test and interpret their physical plausibility. Studies in explainable materials modeling show that transparent model architectures, such as attention mechanisms revealing atomic contributions or saliency mapping over crystal graphs, can perform scientific structural reasoning from numerical output. 150. Such methods are critical for enabling trust calibration, where researchers learn not to over-trusting and under-trusting 49.
@@
LLMs further facilitate the interaction between AI and humans by introducing natural-language interfaces into the workflows. LLMs that are fine-tuned on domain-specific training resources such as experimental protocols and computational reports can effectively parse, summarize, and contextualize scientific results. Recent systems like MatGPT 382 are capable of solving materials language processing tasks by conducting workflows, querying crystallographic databases, and helping experimental setups by giving suggestions in natural language 77. By integrating multimodal reasoning that combines textual, numerical, and structural embeddings, LLMs can also serve as adaptive documentation and knowledge retrieval systems. These systems can evolve with laboratory data streams. In this manner, human–AI collaboration is expansive, from mere task execution to the whole scientific lifecycle: from idea creation to experiment construction and execution, validation, interpretation, and publication 228.
@@
AI systems are becoming more and more autonomous, which makes human scientists become less directly involved in low-level decision-making. This will reduce the participation of human agency and makes co-creation more difficult in some scenarios. More and more AI systems like ChemOS 304 can enhance the planning and execution of experiments in self-driving labs. In this case, human scientists should foucs more on higher-level management, ethical oversight, supervision, and conceptual innovation345. Interface design further contributes to more effective collaboration, which involves designing interpretable feedback loops between algorithmic reasoning and human comprehension 74. For instance, prospective interfaces can prevent confusion before it happens. An explainable user interface allows users to interactively explore the data and model sensitivity so they can predict what and why the AI will do. The AI decision can construct AI results as structured hypotheses, inviting human assessment to decide whether accept or deny them. These hybrid decision-making architectures align with the principles of human-in-the-loop AI, where automation and human expertise together produce more reliable outcomes than either could achieve alone 322.
@@
A second new concept at this stage is scientific collaboration networks, which were originally developed to analyze collaborations among human researchers 31 and has since broadened to include hybrid collaborations between humans and AI agents 125. In traditional scientific collaboration networks, nodes are scientists, and edges connect them if the scientists co-authored at least one paper jointly. Networks were constructed independently for separate fields for comparing the disciplinary patterns. By contrast, nodes in human-AI collaboration networks are specialized agents (software modules) along with human supervision. The framing implicitly favors human-agent teaming: humans define goals, agents create and refine hypotheses, and humans evaluate, execute, and validate results. This structure acts much like a human research team. Different members have different roles that work well together, and they talk to each other through shared interfaces to plan their work.
@@
From an epistemological point of view, the collaboration of human–AI technology signifies a paradigm shift in scientific conduct. Rather than knowledge being created solely through human deliberation, traditional science cycle by hypothesis–experiment–analysis is bounded by human thought and speed of human communication 205. Instead, co-creative workflows exist at the crossroads between symbolic reasoning on the one hand and data-driven discovery, turning this cycle into a process of feedback loop through AI 205. Transforming experimental data into organized form representations understandable to humans and machines, these systems collectively share an epistemic language that brings interpretability and reproducibility 88.
@@
Collaborative frameworks also need to address issues of attribution, authorship, and bias in the future in order to address the trust and accountability issues 261, 265. Designing new architectures and evaluation frameworks for trustworthy collaboration is becoming essential. AutoLabs 275 introduces a critic agent specifically designed to double-check the safety and feasibility, while GIFTERS 17 tries to evaluate the AI part along dimensions such as generalizable, interpretable, fair, transparent, explainable, robust, and stable trustworthiness principles. The scientific community is beginning to deal with questions like: Who owns AI-generated hypotheses? How can we peer review AI-based discoveries? Efforts such as transparent model reporting, tracking data sources, clear documentation, and continuous monitoring are essential for keeping the scientific integrity of AI-assisted research 45.Furthermore, applying fairness, transparency, and accountability principles into materials workflows is crucial for the responsible adoption of AI in both industry and academia 227.